# Inertial Confinement Fusion Forecasting via LLMs

### v2

@@ 5 Conclusion
In this section, we provide a comprehensive discussion and analysis of the confidence scores associated with the Fusion-LLM. As elaborated in §2.2.3, our confidence scores offer per-step evaluations, thereby aiding physicists to gain deeper insights into the reliability across various segments of predictions. This functionality is particularly vital for understanding the model’s performance dynamics within specific contexts of its predictive output.
@@
Such an analysis underscores the importance of confidence scores as a diagnostic tool, highlighting intervals where the model’s predictions are potentially less reliable. By mapping these confidence scores to the corresponding prediction errors, physicists can identify specific phases within the prediction and temporal sequence where the model’s forecasting should be interpreted with caution. This capability not only enhances the trustworthiness of the Fusion-LLM but also provides critical feedback for further refinement of the model in the future research.
@@
The introduction of Fusion-LLM represents a significant advancement in integrating LLMs with classical reservoir computing paradigms to enhance predictive capabilities in Inertial Confinement Fusion. This novel approach not only meets but exceeds several existing state-of-the-art models in performance benchmarks. From a societal perspective, the implications of Fusion-LLM are profoundly beneficial, as our approach provides a valuable tool for advancing our understanding and capabilities in harnessing fusion energy — a potential key to long-term sustainable energy solutions. However, it is imperative to acknowledge and critically assess the potential drawbacks associated with this technology. Similar to other predictive models, Fusion-LLM faces challenges when dealing with out-of-distribution data or scenarios that have not been previously encountered. This limitation underscores the need for ongoing research and refinement, particularly in its application to real-world ICF scenarios where unpredictable behaviors might emerge. Therefore, while the model demonstrates promising applications, its deployment in practical settings must be approached with caution, ensuring continuous evaluation and adaptation to maintain reliability and safety in its predictive assertions.
Appendix S7 Ethical Safeguards
Appendix S8 Reproducibility
Fusion-LLM is implemented in PyTorch [52]. Experiments are conducted on two NVIDIA A100-40GB GPUs. To guarantee reproducibility, our full implementation shall be publicly released upon paper acceptance.
@@
All the methods we used for comparison are publicly available for academic usage. PIC Simulation is implemented based on the reproducing by osiris-code/osiris with AGPL-3.0 license. We use huggingface/transformers for the implementations of Autoformer [86], Llama 2 [79] and Llama 3 [4] with Apache-2.0 license and Llama 2/3 Community License Agreement. We used the official repositories DAMO-DI-ML/NeurIPS2023-One-Fits-All (GPT4TS [94]), KimMeen/Time-LLM [31], rubenohana/Reservoir-computing-kernels (RCRK [17]), CsnowyLstar/HoGRC [43] and quantinfo/ng-rc-paper-code (NGRC [20]) for our comparison experiments, where Time-LLM [31] is licensed under Apache-2.0, HoGRC [43] and NGRC [20] are licensed under MIT, the rest did not mention their licenses.

### v3

@@ 7 Reproducibility
In this section, we provide a comprehensive discussion and analysis of the confidence scores associated with the LPI-LLM. As elaborated in §2.2.3, our confidence scores offer per-step evaluations, thereby aiding physicists to gain deeper insights into the reliability across various segments of predictions. This functionality is particularly vital for understanding the model’s performance dynamics within specific contexts of its predictive output.
@@
Such an analysis underscores the importance of confidence scores as a diagnostic tool, highlighting intervals where the model’s predictions are potentially less reliable. By mapping these confidence scores to the corresponding prediction errors, physicists can identify specific phases within the prediction and temporal sequence where the model’s forecasting should be interpreted with caution. This capability not only enhances the trustworthiness of the LPI-LLM but also provides critical feedback for further refinement of the model in the future research.
@@
The introduction of LPI-LLM represents a significant advancement in integrating LLMs with classical reservoir computing paradigms to enhance predictive capabilities in Inertial Confinement Fusion. This novel approach not only meets but exceeds several existing state-of-the-art models in performance benchmarks. From a societal perspective, the implications of LPI-LLM are profoundly beneficial, as our approach provides a valuable tool for advancing our understanding and capabilities in harnessing fusion energy — a potential key to long-term sustainable energy solutions. However, it is imperative to acknowledge and critically assess the potential drawbacks associated with this technology. Similar to other predictive models, LPI-LLM faces challenges when dealing with out-of-distribution data or scenarios that have not been previously encountered. This limitation underscores the need for ongoing research and refinement, particularly in its application to real-world ICF scenarios where unpredictable behaviors might emerge. Therefore, while the model demonstrates promising applications, its deployment in practical settings must be approached with caution, ensuring continuous evaluation and adaptation to maintain reliability and safety in its prediction.
@@
All the methods we used for comparison are publicly available for academic usage. PIC Simulation is implemented based on the reproducing by osiris-code/osiris with AGPL-3.0 license. We use huggingface/transformers for the implementations of Autoformer (Wu et al., 2021) under Apache-2.0, Llama 2 (Touvron et al., 2023b) under Llama 2 Community License, Llama 3 (AI, 2024) under Llama 3 Community License, Gemma 2 (Team, 2024) under Gemma Terms of Use and OLMo (Groeneveld et al., 2024) under Apache-2.0. We used the official repositories DAMO-DI-ML/NeurIPS2023-One-Fits-All (GPT4TS (Zhou et al., 2023)), KimMeen/Time-LLM (Jin et al., 2023), rubenohana/Reservoir-computing-kernels (RCRK (Dong et al., 2020)), CsnowyLstar/HoGRC (Li et al., 2024) and quantinfo/ng-rc-paper-code (NGRC (Gauthier et al., 2021)) for our comparison experiments, where Time-LLM (Jin et al., 2023) is licensed under Apache-2.0, HoGRC (Li et al., 2024) and NGRC (Gauthier et al., 2021) are licensed under MIT.